In a triangle ABC, the angle at B is 155 degrees, the length of side BC is 29, and the length of side AB is 10. To 2 decimal places, what is the length of side AC?

Answer by Cromlix(4381)

Hi there,
Using Cosine Rule:
b^2 = a^2 + c^2 - 2acCosB
b^2 = 29^2 + 10^2 - 2(29)(10)Cos155
b^2 = 1466.659
b = 38.297
b = 38.3 units
Hope this helps :-)